OpenStudy (anonymous):

Find the 2nd, 3rd, and 5th terms of a geometric sequence having 1st term 8 and 4th term 27.

OpenStudy (campbell_st):

if you are finding a term in a geometric series you can use \[T_{n} = ar^{n-1}\] you know the 1st term a = 8 and you will need to frind r, you also know the 4th term 27 so using the formula you have \[27 = 8r^{4-1}\] divide both sides by 8 \[\frac{27}{8} = r^3\] to find r take the cube root of both sides of the equation. When you have r you will be able to find the 2nd and 3rd terms.

OpenStudy (anonymous):

Alright! So I got that r=3/2... but I'm unsure as to where to go from there. Do we do the same thing (isolate r) to the other known term (a=8)?

OpenStudy (campbell_st):

ok... so the 2nd term n = 2 \[T_{2} = 8 \times (\frac{3}{2})^{2 - 1}\] for the 3rd term n = 3 so substitute that and evaluate and the 5th term is n = 5, substitute and evaluate

OpenStudy (anonymous):

Wow... that's so simple now that you've explained it!! Thank you again, so so much!! :)